Find the coordinates of a point on the line 2x+y-4=0 whose distance from 3x+4y-1=0 is 2 units.


PLEASE REPLY SOON.


Thanks.


 

Grade:

1 Answers

vikas askiitian expert
13 years ago

2x + y - 4 = 0

y = 4-2x

general point on this line is

P = (x,4-2x)

 

distance of this point P from line 3x+4y-1 = 0 is 2 units so

3x+4y-1/5 = 2

3x + 4(4-2x) - 1 = 10

x = 1

so general point P = (x,4-2x) = (1,2)

 

this is the required point

approve if u like my ans